OpenStudy (anonymous):

Which value of x makes the following equation true? 3(x - 3) = -2x + 31 12 8 11 9

OpenStudy (anonymous):

You could plug in each of your values to see which one works. Or, you can do it like this to find x: 3(x-3)=-2x+31 3x-9=-2x+31 3x+2x-9=-2x+2x+31 5x-9=31 5x=40 5x/5=40/5 x=8 Check it: 3(8-3)=-2(8)+31 3(5)=-16+31 15=15 we see that it works!

OpenStudy (anonymous):

thank you :3

OpenStudy (anonymous):

Of course :)